### v2

# CDEX Collaboration

Abstract
[…]
preprint: APS/123-QED
I Introduction
There is strong evidence indicating that dark matter (DM) is a major constituent of the Universe Bertone and Tait (2018); Workman et al. (2022); Matarrese et al. (2011); Bertone et al. (2005). Direct detection (DD), indirect detection Ambrosi et al. (2017), and collider Daci et al. (2015); Jing et al. (2023); Chen et al. (2023); Wang (2022) experiments have been conducted for decades to search for DM particles (denoted as “$\chi$”). In DD experiments, the most popular search object is the $\chi$ in the Standard Halo Model (SHM). The velocity of $\chi$ in the SHM follows the Maxwell-Boltzmann distribution with a most probable value of 238 km/s and a cutoff at 544 km/s Smith and Others (2007); Baxter et al. (2021). Time projection chambers (XENON Aprile et al. (2023), LUX-ZEPLIN Aalbers et al. (2023), PandaX Li et al. (2023), DarkSide Agnes et al. (2018)), cryogenic calorimeters (CRESST Abdelhameed et al. (2019), SuperCDMS Amaral et al. (2020), EDELWEISS Arnaud et al. (2020)), charge-coupled devices (SENSEI Barak et al. (2020), DAMIC Aguilar-Arevalo et al. (2019)), and high purity germanium detectors (CoGeNT Aalseth et al. (2013), CDEX Liu et al. (2014); Zhao et al. (2013); Yue et al. (2014); Zhao et al. (2016); Yang et al. (2018); Jiang et al. (2018, 2019); Yang et al. (2019); Wang et al. (2021); Geng et al. (2023); Zhang et al. (2022)) have been operated in searching for $\chi$ by investigating their interaction with target electrons or target nuclei.
In the paradigm of the $\chi$–nucleus ($\chi$$N$) scattering, the best detection sensitivity has been achieved for $m_{\chi}\sim\mathcal{O}$(10 GeV) Aalbers et al. (2023), where $m_{\chi}$ is the mass of $\chi$. As $m_{\chi}$ decreases, little energy transfers to the heavy nucleus in the elastic scattering process, resulting in a rapid decrease in the detection sensitivity. Considering the Migdal effect, the sensitivity in the low $m_{\chi}\sim\mathcal{O}$(100 MeV) range can be significantly improved Liu et al. (2019); Akerib et al. (2019); Angloher et al. (2017); Liu et al. (2022). Further, when $\chi$ are scattered and accelerated to (semi)relativistic velocities by cosmic ray nuclei Cappiello and Beacom (2019); Xu et al. (2022); Bringmann and Pospelov (2019); Cappiello et al. (2019); Ema et al. (2021); Andriamirado et al. (2021); Cui et al. (2022), blazar jet protons Wang et al. (2022), and supernova remnants Cappiello et al. (2023), $\chi$ with $m_{\chi}\lesssim\mathcal{O}$(1 GeV) can be searched.
In another paradigm parallel to $\chi$$N$ scattering, the $\chi$–electron ($\chi$$e$) scattering, the best detection sensitivity has been achieved for SHM $\chi$ with a mass of $m_{\chi}\sim\mathcal{O}$(10 MeV) Zhang et al. (2022) rather than $m_{\chi}\sim\mathcal{O}(m_{e})=\mathcal{O}$(0.5 MeV) because of the reduced $\chi$ kinetic energy. Similar to the $\chi-N$ scenario, $\chi$ can be accelerated by cosmic ray electrons Cappiello and Beacom (2019); Cappiello et al. (2019); Ema et al. (2019), cosmic ray neutrinos Yin (2019), stellar neutrinos Zhang (2022); Jho et al. (2021), diffuse supernova neutrinos Das and Sen (2021); Ghosh et al. (2022), stellar electrons An et al. (2018, 2021); Emken (2022); Emken et al. (2018), and blazar jet electrons Bhowmick et al. (2023); Granelli et al. (2022), a much lower $m_{\chi}$ can be searched.
In addition to $\chi$ and target elastic scattering, some branching models, such as neutral current absorption Dror et al. (2020); Dai et al. (2022) and effective electromagnetic interaction Ning et al. (2023); Bai (2023), have increasingly gained attention in DD experiments. Moreover, cosmic rays bombarding the atmosphere are used for $\chi$ detection Plestid et al. (2020); Alvey et al. (2019); Arguëlles et al. (2022); Flambaum et al. (2023); Yuan (2023). In this work, we investigate a novel $\chi$ source in DD experiments: evaporating primordial black holes (PBHs).
PBH is a possible candidate for DM DeRocco and Graham (2019); Green and Kavanagh (2021); Belotsky et al. (2014); Dasgupta et al. (2020). Gravitational collapses due to density fluctuations may have resulted in the formation of PBHs soon after the inflationary epoch of the early Universe Zel’dovich and Novikov (1967); Harrison (1970); Khlopov (2010). According to general relativity and quantum field theories, when a PBH evaporates, it emits particles, which is well known as Hawking radiation Hawking (1974). The $\chi$ can also be characterized as Hawking radiation Morrison et al. (2019); Baldes et al. (2020); Bernal and Zapata (2021a, b, c); Arbey and Auffinger (2021); Gondolo et al. (2020); Masina (2021); Cheek et al. (2022a, b). An evaporating PBH is a novel source of $\chi$ Calabrese et al. (2022a, b); Li and Liao (2022); Zhang et al. (2023) or other new particles Li and Zhang (2022); Agashe et al. (2023); Jho et al. (2022) beyond the standard model (SM). For convenience, the scenario of “a PBH evaporating $\chi$” is abbreviated as “PBHeDM”. PBHeDM is model-independent only when considering the gravitational interactions between the dark sector and SM Friedlander et al. (2023); model-dependence comes in when numerical values are derived because of the calculation models.
[…]
II The $\chi$ emitted from evaporating PBHs
In 1974, Hawking proposed that evaporating black holes emit matter in the form of thermal radiation Hawking (1974). The relation between the Hawking temperature ($T_{PBH}$) and mass ($M_{PBH}$) of a Schwarzschild black hole is given by the Zeroth law of thermodynamics for black holes based on thermodynamics and the gravitational theory Page (1976, 1977); MacGibbon and Webber (1990), as described in Eq. 1. We investigate black holes with a mass of $M_{PBH}\sim\mathcal{O}$(1016 g). The mass of these black holes is very small compared to the mass of the Sun; consequently, they are considered primordial rather than cosmological black holes. Thus, we directly refer to these black holes as “PBHs”.
$k_{B}T_{PBH}=\frac{\hbar c^{3}}{8\pi GM_{PBH}},$ (1)
where $\hbar$ is the reduced Planck constant, $c$ is the speed of light in vacuum, $G$ is the Newtonian constant of gravitation, and $k_{B}$ is the Boltzmann constant.
A particle $\xi$ emitted from an evaporating PBH can be not only SM particles (such as a photon, a neutrino, or an electron) but also dark radiation. Its emission rate is described in Eq. 2 and can be computed using the publicly available $\tt Blackhawk$-$\tt v2.1$ code Arbey and Auffinger (2019, 2021).
$\frac{{\rm d}^{2}N_{\xi}}{{\rm d}E_{\xi}{\rm d}t}=\frac{g_{\xi}}{2\pi}\frac{\Gamma_{\xi}(E_{\xi},M_{PBH},x)}{exp(E_{\xi}/k_{B}T_{PBH})-(-1)^{2s_{\xi}}},$ (2)
where $E_{\xi}$ is the evaporated $\xi$ energy, $x$ is a set of secondary parameters for the PBH metrics, and $\Gamma_{\xi}$ is the graybody factor, which represents the probability that a particle generated at the PBH event horizon will escape to infinity in space.
In this work, $\xi=\chi$. Here, we consider the simple case of Dirac fermions ($s_{\chi}=1/2$) with four degrees of freedom ($g_{\chi}=4$) for a chargeless and spinless PBH with a monochromatic distribution $M_{PBH}$ as an example.
III The $\chi$ flux reaching the Earth
The $\chi$ flux emitted from an evaporated PBH in the Milky Way (MW) galaxy and reaches the Earth can be described as follows:
$\frac{{\rm d}^{2}\phi_{\chi}^{MW}}{{\rm d}T_{\chi}{\rm d}\Omega}=\frac{f_{PBH}}{4\pi M_{PBH}}\int\frac{{\rm d}\Omega_{s}}{4\pi}\int_{\rm l.o.s.}\rho_{MW}^{\rm NFW}[r(s,\phi)]\frac{d^{2}N_{\chi}}{dT_{\chi}dt}{\rm d}s,$ (3)
where the kinetic energy of $\chi$ is $T_{\chi}=E_{\chi}-m_{\chi}$, $\rho_{MW}^{\rm NFW}$ is the DM density of the Milky Way halo in the Navarro-Frenk-White (NFW) profile Navarro et al. (1997) with a local density $\rho_{\odot}$ = 0.4 GeV/cm3, $r(s,\phi)$ is the galactocentric distance (the solar distance from the galactic center), $s$ is the line-of-sight (l.o.s.) distance to the PBH, $\phi$ is the angle between the two directions, and $f_{PBH}$ is defined as the fraction of DM composed of PBHs ($f_{PBH}=\rho_{PBH}/\rho_{DM}$).
The constraint results in the scenario of galactic $\gamma$-ray have a slight difference Coogan et al. (2021) between the Einasto Einasto (1965) and NFW Navarro et al. (1997) profiles. This difference is even smaller after combining the Hawking radiation fluxes from the extragalactic (EG) PBHs.
EDGES 21 cm Clark et al. (2018); Mittal et al. (2022) and COMPTEL Coogan et al. (2021) have imposed the most stringent constraints on $f_{PBH}$ for PBHs with masses ranging from $1\times 10^{15}$ to $7\times 10^{16}$ g. The evaporation of PBH heats the original gas and weakens the signal of the 21 cm absorption line. Clark et al. Clark et al. (2018) derive the limit on $f_{PBH}$ from EDGES data. Coogan et al. Coogan et al. (2021)derive the limit on $f_{PBH}$ from COMPTEL data based on the $\gamma$ evaporation from the galactic PBHs. The $f_{PBH}$ values selected in this work are listed in Table 1. A cutoff was applied to $M_{PBH}=1\times 10^{15}$ g since PBHs with $M_{PBH}\lesssim\mathcal{O}(10^{14})$ g evaporate violently and decay away; a cutoff was applied to $M_{PBH}=7\times 10^{16}$ g since the existing experimental constraints on $f_{PBH}$ for large $M_{PBH}$ are close to or larger than 0.1. In addition, there are no valid observational limits on $f_{PBH}$ for $M_{PBH}\sim\mathcal{O}$(1017–1023) g Carr et al. (2021); Green and Kavanagh (2021).
Table 1: Selected $f_{PBH}$ values for different $M_{PBH}$s, based on the latest constraints obtained from Refs. Clark et al. (2018); Coogan et al. (2021).
   $M_{PBH}$ (g) $f_{PBH}$ $M_{PBH}$ (g) $f_{PBH}$
   $1\times 10^{15}$ $1.6\times 10^{-8}$ $9\times 10^{15}$ $5.2\times 10^{-5}$
   $2\times 10^{15}$ $1.8\times 10^{-7}$ $1\times 10^{16}$ $7.0\times 10^{-5}$
   $3\times 10^{15}$ $9.8\times 10^{-7}$ $2\times 10^{16}$ $1.0\times 10^{-3}$
   $4\times 10^{15}$ $3.5\times 10^{-6}$ $3\times 10^{16}$ $3.8\times 10^{-3}$
   $5\times 10^{15}$ $8.6\times 10^{-6}$ $4\times 10^{16}$ $7.6\times 10^{-3}$
   $6\times 10^{15}$ $1.4\times 10^{-5}$ $5\times 10^{16}$ $1.4\times 10^{-2}$
   $7\times 10^{15}$ $2.0\times 10^{-5}$ $6\times 10^{16}$ $2.2\times 10^{-2}$
   $8\times 10^{15}$ $3.5\times 10^{-5}$ $7\times 10^{16}$ $3.4\times 10^{-2}$
The effect of the cosmological red-shift $z(t)$ on the particle kinetic energy should also be taken into account for the EG component:
$\frac{{\rm d}^{2}\phi_{\chi}^{EG}}{{\rm d}T_{\chi}{\rm d}\Omega}=\frac{f_{PBH}\rho_{DM}}{4\pi M_{PBH}}\int^{t_{U}}_{t_{0}}[1+z(t)]\frac{d^{2}N_{\chi}}{dT_{\chi}dt}{\rm d}t,$ (4)
where $\rho_{DM}$ = 2.35$\times$10-30 g/cm3 is the average DM density of the Universe at the current epoch Aghanim, N. et al. (2020), $t_{0}=10^{11}$ s is the Universe time of the matter-radiation equality, and $t_{U}=1.38\times 10^{10}$ yr is the Universe age; the $\chi$ energy at the source $E_{S}$ and the $\chi$ energy at the Earth’s frame $E_{\chi}$ obey $1+z(t)=[(E^{2}_{S}-m_{\chi}^{2})/(E^{2}_{\chi}-m_{\chi}^{2})]^{1/2}$.
Figure 1: Fluxes of $\chi$ with a mass $m_{\chi}$ = 0.1 keV reaching the Earth from PBHs with masses $M_{PBH}$ = $5\times 10^{16}$ g. The components from galactic (MW) and extragalactic (EG) PBHs are shown by the blue dotted-dashed line and the red dashed line, respectively. $f_{PBH}$s (the fractions of DM composed of PBHs) selected for our calculations are listed in Table 1.
In Eq. 5, the $\chi$ flux from a PBH to the full sky reaching the Earth has two components: the flux from MW PBHs and the flux from EG PBHs.
$\frac{{\rm d}^{2}\phi_{\chi}}{{\rm d}T_{\chi}{\rm d}\Omega}=\frac{{\rm d}^{2}\phi_{\chi}^{MW}}{{\rm d}T_{\chi}{\rm d}\Omega}+\frac{{\rm d}^{2}\phi_{\chi}^{EG}}{{\rm d}T_{\chi}{\rm d}\Omega},$ (5)
Figure 1 shows the flux of $\chi$ with a mass $m_{\chi}$ = 0.1 keV reaching the Earth from galactic and extragalactic PBHs with masses $M_{PBH}$ = $5\times 10^{16}$ g. Considering the EG component, the cosmological red-shift significantly affects $T_{\chi}$ and makes a significant contribution to the flux in the low $T_{\chi}$.
IV Earth shielding
When $\chi$ arrive at the Earth, they must travel through the Jinping Mountain before reaching the detector in CJPL. Similar to the target electrons in a detector, the electrons in the rocks may also scatter with $\chi$. For a large scattering cross section with a mean free path in the rocks of less than a few km, $\chi$ will scatter with the electrons in the rocks and lose kinetic energy before reaching the detectors at underground locations. This phenomenon is called “Earth attenuation” or “Earth shielding effect”. We employ a ballistic-trajectory approximation based on Refs. Bringmann and Pospelov (2019); Calabrese et al. (2022a); Li and Liao (2022).
The kinetic energy of $\chi$ is reduced by ${\rm d}T$ per length ${\rm d}l$ due to its isotropic elastic scattering caused by the electrons in the medium. After traveling a distance $d$, the kinetic energy of $\chi$ is reduced from $T_{0}$ to $T_{d}$, and the flux of $\chi$ changes from $\frac{{\rm d}^{2}\phi_{\chi}}{{\rm d}T_{0}{\rm d}\Omega}$ to $\frac{{\rm d}^{2}\phi_{\chi}^{d}}{{\rm d}T_{d}{\rm d}\Omega}$, where
$\frac{{\rm d}^{2}\phi_{\chi}^{d}}{{\rm d}T_{d}{\rm d}\Omega}\approx\frac{4m_{\chi}^{2}e^{\tau}}{(2m_{\chi}+T_{d}-T_{d}e^{\tau})^{2}}\left(\frac{{\rm d}^{2}\phi_{\chi}}{{\rm d}T_{0}{\rm d}\Omega}\right),$ (6)
with
$T_{d}(T_{0})=\frac{2m_{\chi}T_{0}}{2m_{\chi}e^{\tau}-T_{0}(1+e^{\tau})},$ (7)
where $\tau=d/l$, $d$ = 2400 m is the rock overburden depth of the CJPL Cheng et al. (2017), and $l$ is the interaction length given as follows:
$l=\left[\bar{n}_{e}\sigma_{\chi e}\frac{2m_{e}m_{\chi}}{(m_{e}+m_{\chi})^{2}}\right]^{-1},$ (8)
where $m_{e}$ is the electron mass and $\sigma_{\chi e}$ is the $\chi$–electron elastic-scattering cross section. The average number density of the electrons in the rocks ($\bar{n}_{e}=8.1\times 10^{23}$ cm-3) is computed from the density and the elemental composition of the rocks, as described in Ref. Liu et al. (2022).
V Expected spectra in the detector
When $\chi$ arrives at the detector, there is a probability of elastic scattering and leaving some recoil energy of $E_{r}$. If $E_{r}$ is higher than the binding energy of a shell electron, the event may be recorded by the detector. For $\chi$, the number of target electrons is $\left(N_{Ge}Z^{Ge}_{eff}\sigma_{\chi e}\right)$, where $N_{Ge}$ = 8.30$\times$1021 g-1 is the number density of germanium atoms and $Z^{Ge}_{eff}$ is the effective electron number of germanium listed in Table 2 GeE (2023); Thompson et al. (2009). For instance, when $E_{r}=$ 1 keV in a scattering, then 22 out of 32 shell electrons of the target atom are considered target candidates by $\chi$, and the effective electron number is 22. Considering once again that the scattering is isotropic, the differential event rate of the $\chi$-electron scattering in the detector is given in Eq. 9 Corona et al. (2023); Li and Liao (2022).
$\frac{{\rm d}R_{r}}{{\rm d}E_{r}}(E_{r})=N_{Ge}Z^{Ge}_{eff}\sigma_{\chi e}\int\frac{{\rm d}^{2}\phi_{\chi}^{d}}{{\rm d}T_{d}{\rm d}\Omega}\frac{\Theta(E_{r}^{max}-E_{r})}{E_{r}^{max}}{\rm d}T_{d}{\rm d}\Omega,$ (9)
where $\Theta$ is the Heaviside function and $E_{r}^{max}$ is the maximum value of the recoil energy $E_{r}$ limited by
$E_{r}^{max}=\frac{T_{d}^{2}+2m_{\chi}T_{d}}{T_{d}+(m_{\chi}+m_{e})^{2}/(2m_{e})}.$ (10)
When estimating the average scattering cross section $\bar{\sigma}_{\chi e}$ according to Ref. Calabrese et al. (2022b), calculation is difficult as germanium crystal has a band structure and Ge atoms cannot be considered isolated Griffin et al. (2021); Zhang et al. (2022). Moreover, most $\chi$ are relativistic particles. Therefore, a more elaborate approach should be considered in future calculations.
By convoluting the energy resolution obtained from the calibration, the recoil event rate $\frac{{\rm d}R_{r}}{{\rm d}E_{r}}$ is converted into the expected event rate $\frac{{\rm d}R_{e}}{{\rm d}E_{e}}$ recorded by the CDEX-10 detector. The energy resolution in the CDEX-10 experiment is $\sigma_{Res}(E)=35.8+16.6\cdot E^{1/2}$ eV, where $E$ is in keV units Jiang et al. (2019); Zhang et al. (2022). The spectra before (blue solid line) and after (red dashed line) considering the energy resolution are shown in Fig. 2. There are some step-like fine spectral structures with the binding energy of Ge atom, which are retained with the excellent energy resolution of the CDEX-10 detector, although somewhat smoothed.
Table 2: Effective electron number of germanium $Z^{Ge}_{eff}$ GeE (2023); Thompson et al. (2009).
    $Z^{Ge}_{eff}$   $E_{r}$ (keV)
32 1.11e1 $<E_{r}$
30 1.41e0 $<E_{r}\leq$ 1.11e1
28 1.25e0 $<E_{r}\leq$ 1.41e0
24 1.22e0 $<E_{r}\leq$ 1.25e0
22 1.80e-1$<E_{r}\leq$ 1.22e0
20 1.25e-1$<E_{r}\leq$ 1.80e-1
16 1.21e-1$<E_{r}\leq$ 1.25e-1
14 2.98e-2$<E_{r}\leq$ 1.21e-1
10 2.92e-2$<E_{r}\leq$ 2.98e-2
4 7.20e-3$<E_{r}\leq$ 2.92e-2
0 $E_{r}\leq$ 7.20e-3
Figure 2: Recoil energy (blue solid line) and expected energy (red dashed line) event rates for $M_{PBH}=5\times 10^{16}$ g, $m_{\chi}=0.1$ keV, and $\sigma_{\chi e}=1.3\times 10^{-29}$ cm2. The energy resolution in the CDEX-10 experiment is $\sigma_{Res}(E)=35.8+16.6\cdot E^{1/2}$ eV, where $E$ is in keV units Jiang et al. (2019); Zhang et al. (2022). The fine black dashed line indicates the 160-eV analysis threshold.
VI Results
[…]
1) The zero peak (random trigger events), the 8.98 keV peak (65Zn), the 10.37 keV peak (68Ge) were used to calibrate the energy in the low-energy region.
2) The events with anomalous electronic noise profiles were rejected in the pedestal cut.
3) Based on the linear relationship between the integral value and the maximum value of the physical event pulse, the physicals signal can be discriminated from the electronic noises near the energy threshold.
4) Events depositing energy in the surface layer of the detector generate a slow rising pulse and an incomplete charge collection. The energy measured in these events are low and should be removed.
The measured spectrum is shown by the black points with error bars in Fig. 3. Then, a minimum-$\chi^{2}$ analysis, as described in Eq. 11, is performed. The expected event rate $\frac{{\rm d}R_{e}}{{\rm d}E_{e}}$, denoted by $S$, is a function of $\left(M_{PBH},f_{PBH},m_{\chi},\sigma_{\chi e}\right)$. The background $BKG$ depends on $\left(a,b,I_{K},\sigma_{K},\sigma_{L},\sigma_{M}\right)$ following Eq. 12. The best estimate of $BKG$ can be evaluated by scanning this parameter space in the minimum-$\chi^{2}$ analysis, from which $S$ can be derived. The components of the best fit result are plotted in Fig. 3.
[…]
Table 3: Identified X-ray peaks in the CDEX-10 measured spectrum.
   $E$ (keV) Nuclide Shell $E$ (keV) Nuclide Shell
   0.14 68Ga $M$ 1.30 68Ge $L$
   0.16 68Ge $M$ 4.97 49V $K$
   0.56 49V $L$ 5.99 54Mn $K$
   0.70 54Mn $L$ 6.54 55Fe $K$
   0.77 55Fe $L$ 8.98 65Zn $K$
   1.10 65Zn $L$ 9.66 68Ga $K$
   1.19 68Ga $L$ 10.4 68Ge $K$
$\chi^{2}$ $\left(M_{PBH},f_{PBH},m_{\chi},\sigma_{\chi e}\right)=$ (11)
$\sum_{i}\frac{\left[n_{i}-BKG_{i}-S_{i}\left(M_{PBH},f_{PBH},m_{\chi},\sigma_{\chi e}\right)\right]^{2}}{\Delta^{2}_{i}},$
where $n_{i}$ and $\Delta_{i}$ are the measured data and the standard deviation of the statistical and systematic components at the $i$-th energy bin, respectively.
$BKG=\left(a\cdot E+b\right)$ $+$ $\sum_{M}\frac{I_{M}}{\sqrt{2\pi}\sigma_{M}}exp{\left[-\frac{(E-E_{M})^{2}}{2\sigma_{M}^{2}}\right]}$ (12)
$+$ $\sum_{L}\frac{I_{L}}{\sqrt{2\pi}\sigma_{L}}exp{\left[-\frac{(E-E_{L})^{2}}{2\sigma_{L}^{2}}\right]}$
$+$ $\sum_{K}\frac{I_{K}}{\sqrt{2\pi}\sigma_{K}}exp{\left[-\frac{(E-E_{K})^{2}}{2\sigma_{K}^{2}}\right]},$
where $\left(a\cdot E+b\right)$ is the linear platform; $I_{M}$s, $I_{L}$s, and $I_{K}$s are the intensities of the $M$, $L$, and $K$-shell X ray peaks, respectively; $\sigma_{M}$s, $\sigma_{L}$s, and $\sigma_{K}$s are the energy resolutions of the $M$, $L$, and $K$-shell X-ray peaks, respectively. Referring to our previous studies Xu et al. (2022); Dai et al. (2022), fourteen X-ray peaks were identified and listed in Table 3. $I_{K}$s are used to constrain $I_{M}$s and $I_{L}$s with known $K$/$L$ and $K$/$M$ ratios Bahcall (1963); Jiang et al. (2018).
Figure 4: Exclusion regions ($m_{\chi}$, $\sigma_{\chi e}$) in PBHeDM with a 90% CL for $M_{PBH}$ = (1, 10, 70)×$\times 10^{15}$ g, and $f_{PBH}=1.6\times 10^{-8},7.0\times 10^{-5},3.4\times 10^{-10}$ respectively as listed in Table 1. For comparison, the limit on the cold $\chi$ in the SHM Barak et al. (2020); Aguilar-Arevalo et al. (2019); Arnaud et al. (2020); Agnese et al. (2019); Agnes et al. (2018); Aprile et al. (2019); Cheng et al. (2021) is also shown in dashed gray shadow. The results obtained from the phenomenological interpretations of Li et al. Li and Liao (2022) for the Super-Kamiokande and XENON-1T data for $M_{PBH}=10^{15}$ g and $f_{PBH}=3.9\times 10^{-7}$ are plot respectively in red and blue. For visual clarity, the results from Calabrese et al. Calabrese et al. (2022b) are omitted from the figure.
[…]
In the new concept of PBHeDM, the parameter space of ($m_{\chi}$, $\sigma_{\chi e}$) can be effectively searched in the low $m_{\chi}$. The results obtained from the phenomenological interpretations of Li et al. Li and Liao (2022) for the Super-Kamiokande and XENON-1T data for $M_{PBH}=10^{15}$ g and $f_{PBH}=3.9\times 10^{-7}$ are plot respectively in red and blue. In Calabrese et al.’s work Calabrese et al. (2022b), the atom ionization with an outgoing nonrelativistic free electron caused by the $\chi$$e$ scattering is treated by the theory of nondenormalized effective fields, and then the lower limits on the mean cross section $\bar{\sigma}_{\chi e}$ are derived from the Super-Kamiokande and XENON-1T data. For visual clarity, the results are omitted from the figure. CDEX-10’s radiation background is about three orders of magnitude higher than XENON-1T’s. But CDEX-10’s energy threshold (0.16 keV) is lower than XENON-1T’s (1 keV). Thus, CDEX-10 excludes the space above them.
As a supplement, the projection of ($M_{PBH}$, $f_{PBH}$, $m_{\chi}$ = 0.1 keV, $\sigma_{\chi e}$) on ($M_{PBH}$, $\sigma_{\chi e}$) is drawn in Fig. 5 numerically but not physically because both the upper- and lower-bound lines of the exclusion regions ($m_{\chi}$, $\sigma_{\chi e}$) for $m_{\chi}$ $\lesssim$ 0.1 keV are flat.
Figure 5: Numerical exclusion regions ($M_{PBH}$, $\sigma_{\chi e}$) with a 90% CL for $m_{\chi}$ $\lesssim$ 0.1 keV. The lower and upper bounds on $\sigma_{\chi e}$ are shown by the blue and red solid lines, respectively. The gray-shaded parameter space in between is probed and excluded.
VII Discussion
To calculate the physical ($m_{\chi}$, $\sigma_{\chi e}$) and numerical ($M_{PBH}$, $\sigma_{\chi e}$) results, the $f_{PBH}\left(M_{PBH}\right)$ values listed in Table 1 used as inputs in Eq. 3 and 4. The flux reaching the Earth of $\frac{{\rm d}^{2}\phi_{\chi}}{{\rm d}T_{\chi}{\rm d}\Omega}$ in Eq. 5 is proportional to the $f_{PBH}$. The red line in Fig. 5 has a step-wise structure at $M_{PBH}\simeq 1.5\times 10^{16}$ g because it is required the stronger Earth attenuation to reduce the flux reaching the detector. However, the effect of Earth shielding on the lower limit is not significant and the $\frac{{\rm d}^{2}\phi_{\chi}^{d}}{{\rm d}T_{d}{\rm d}\Omega}$ in Eq. 6 is approximately equal to the $\frac{{\rm d}^{2}\phi_{\chi}}{{\rm d}T_{\chi}{\rm d}\Omega}$ in Eq. 5. The blue line does not vary as rapidly as the red line because the expected spectrum exhibits some fine shapes.
There remains a degenercy in $\sigma_{\chi e}(m_{\chi})$ and $f_{PBH}(M_{PBH})$ from this work. If ($m_{\chi},\sigma_{\chi e}$) can be determined in the future in other scenarios, such as SHM, we can search for evaporating PBHs through $\chi$ as particles in the SM such as $\gamma$ photons. Here, $\sigma_{\chi e}$ in turn becomes the input value and $f_{PBH}$ becomes the experimental output limit in DD experiments. For $m_{\chi}\lesssim 0.1$ keV, upper bounds with a 90% CL on $f_{PBH}\left(M_{PBH}\right)$ for different $\sigma_{\chi e}$s from CDEX-10 are plotted in lines in Fig. 6.
Figure 6: Upper bounds with a 90% CL on $f_{PBH}\left(M_{PBH}\right)$ for different $\sigma_{\chi e}$s ($m_{\chi}$ $\lesssim$ 0.1 keV). The numbers $5\times 10^{-27}$, $1\times 10^{-27}$, and so on, shown in the legend are in cm2 and correspond to $\sigma_{\chi e}$s. Some strong existing bounds obtained from the extragalactic $\gamma$-ray background (EGR) Carr et al. (2010), EDGES 21 cm Clark et al. (2018), Voyager-1 Boudaud and Cirelli (2019), and COMPTEL Coogan et al. (2021) observations are shown; the strongest bounds marked with black circles correspond to $f_{PBH}$ listed in Table 1 and were selected as inputs in Eqs. 4 and 5. The results obtained from the phenomenological interpretations of Calabrese et al. Calabrese et al. (2022b) and Li et al. Li and Liao (2022) for the Super-Kamiokande and XENON-1T data are omitted from the figure for visual clarity.
All of the $f_{PBH}$s (including those from COMPTEL from other experiments) increase as $M_{PBH}$ increase because of the decreased Hawking temperature. For larger $M_{PBH}$, a lower energy threshold in the analysis is required. Due to the effect of Earth shielding, $f_{PBH}$ decreases with the decrease in $\sigma_{\chi e}$ between $\mathcal{O}(5\times 10^{-27})$ and $\mathcal{O}(1\times 10^{-28})$ cm2 in Fig. 6. Therefore, an accurate Earth shielding model rather than an approximation is required. This remains an open research area that will be the theme of our future efforts. Between $\mathcal{O}(1\times 10^{-28})$ and $\mathcal{O}(1\times 10^{-29})$ cm2, $f_{PBH}$ increases with the decrease in $\sigma_{\chi e}$.
It is worth noting that in Fig. 6, the CDEX-10 curve rises more slowly than the black dots (which is the current best $f_{PBH}$ limit obtained from cosmological observations Carr et al. (2010); Clark et al. (2018); Boudaud and Cirelli (2019); Coogan et al. (2021)). The XENON-1T’s results shown in Fig. 3 in Ref. Calabrese et al. (2022b) and in Fig. 4 in Ref. Li and Zhang (2022) exhibit the same trend. For example, the curve of ($m_{\chi}=$ 0.1 keV, $\sigma_{\chi e}=1\times 10^{-29}$ cm2) is out of the shaded part when $M_{PBH}\gtrsim 8\times 10^{16}$ g. This indicates that PBHeDM could be used to search for evaporating PBHs with large $M_{PBH}$s and imposes constraints on $f_{PBH}$s. DD experiments with low radiation background and low energy threshold, such as XENON and CDEX, are expected to fill the gap in ($M_{PBH}$, $f_{PBH}$) for $M_{PBH}\sim\mathcal{O}(10^{18})$ g in Ref. Carr et al. (2021); Green and Kavanagh (2021) if the physical properties of $\chi$ can be clarified in the future.
VIII Summary
PBHeDM is a novel non-thermal DM particle. In the presence of ($M_{PBH}$, $f_{PBH}$) compatible with present bounds, CDEX-10 experiment excludes $\sigma_{\chi e}\sim\mathcal{O}(10^{-28})$ cm2 for $\chi$ with a mass $m_{\chi}\lesssim$ 0.1 keV for $M_{PBH}$ ranging from $1\times 10^{15}$ g to $7\times 10^{16}$ g. This is a parameter space that previous works have not explored. In the future, if $\sigma_{\chi e}\left(m_{\chi}\right)$ is known, then DD experiments will have the potential to give limits on $f_{PBH}\left(M_{PBH}\right)$ for the larger $M_{PBH}$.
Acknowledgements
This work was supported by the National Key Research and Development Program of China (Grants No. 2023YFA1607100 and No. 2022YFA1605000) and the National Natural Science Foundation of China (Grants No. 12322511, No. 12175112, No. 12005111, and No. 11725522). We acknowledge the Center of High performance computing, Tsinghua University for providing the facility support. We would like to thank CJPL and its staff for hosting and supporting the CDEX project. CJPL is jointly operated by Tsinghua University and Yalong River Hydropower Development Company.